Lance Whittingham IV specializes in buying deep discount bonds. These represent bonds that are trading at well below par value. He has his eye on a bond issued by the Leisure Time Corporation. The $1,000 par value bond pays 4 percent annual interest and has 18 years remaining to maturity. The current yield to maturity on similar bonds is 11 percent. Use Appendix B and Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods.

a. What is the current price of the bonds? (Do not round intermediate calculations. Round your final answer to 2 decimal places. Assume interest payments are annual.)

Current price of the bond   

b. By what percent will the price of the bonds increase between now and maturity? (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)

Price increases by %
Answer #1

Price of a bond= 21 (1+r)t + (1+r)7] Coupon + F

Where,

Coupon is the amount of coupon payable each period

r is the rate of yield to maturity

F is the maturity value

T is the period of maturity

By using the above formula, we can answer the questions as below-

  1. Price of the bond= ΠΓ 40 1000 Σι1+011t (1+0.11)18)

Where, t is 1 to 18. Therefore, we should take PVIFA factor of 11% for 18 years and multiply it with coupon amount.

PVIFA for 11% for 18 years (Check PVIFA chart) is 7.701617

Therefore, cash inflows from coupon is = 7.701617×40= 308.0647

And,

1000 *(1+0.11)18]= 152.8222

              Hence, price of the bond= 308.0647+152.8222= 460.89

b) Changes in price of bond during the period till maturity = (maturity value - Current price)/Current price× 100

= (1000-460.89)/460.89× 100

= 116.97%